OpenStudy (anonymous):

What is the equation of a line, in general form, with a slope of -2 and a y-intercept of 8? x + 2y - 8 = 0 2x + y - 8 = 0 2x - y + 8 = 0

OpenStudy (anonymous):

option 2

OpenStudy (anonymous):

Yeah, thanks for helping -.- Please don't give out direct answers.

OpenStudy (britanyashton):

I can explain it to you, if you'd like Nattie.

OpenStudy (anonymous):

Thanks,@britanyashton that would be appreciated, and you can just call me Nat. :)

OpenStudy (britanyashton):

How I would do this- well I remember y=mx+b. the slope is -2 and the b=8 -2x+8=y -y -y -2x-y+8=0 I would multiply each side of it by -1 to cancel out that negative sign. 2x+y-8=0 I haven't done this in forever, but i can almost say that this is right.

OpenStudy (anonymous):

Lol, that makes sense, thanks CX

OpenStudy (britanyashton):

otay. If its wrong, im soo sorry! :) Good luck!

OpenStudy (anonymous):

Lol thanks CX I bet you are right, that looks to be what I would do CX